OpenStudy (anonymous):

How to start on --> x=2 square root of x-1? Help please...

OpenStudy (anonymous):

\[x=2\sqrt{x-1}\]

OpenStudy (gorv):

take square of both sides

OpenStudy (gorv):

\[x^{2}=(2*\sqrt{x-1})^{2}\]

OpenStudy (gorv):

x^2=2^2*(x-1) x^2=4(x-1)

OpenStudy (gorv):

solve it further

OpenStudy (anonymous):

got it thank you!